the sum of the three positive integers is 315. find the numbers?
Answer by Alan3354(69443)

x + y + z = 315
There are a lot of combinations that fit.
1 + 2 + 312 = 315
1 + 3 + 311 = 315
1 + 4 + 310 = 315
5 + 10 + 300 = 315
etc